OpenStudy (linn99123):

NEED HELP FAST WILL GIVE FAN AND MEDAL :D West deposited $15,500 in a savings account earning 2.5% interest. What was the total interest earned at the end of 8 years? A. $310.00 B. $484.38 C. $3,100.00 D. $4,843.75

OpenStudy (the_fizicx99):

You try it boo

OpenStudy (anonymous):

If Kim wanted to create a function that modeled a base of 7 and what exponents were needed to reach specific values, how would she set up her function? f(x) = x^7 f(x) = 7^x f(x) = log7x f(x) = logx7

OpenStudy (linn99123):

can ya help me :o

OpenStudy (anonymous):

ya

OpenStudy (the_fizicx99):

@Jedi3000 ask your questions on your on post. ._. we just did this ;~; 15,500 * 2.5 38,750 * 8 = 310,000

OpenStudy (anonymous):

It is C

OpenStudy (the_fizicx99):

no shi-t lol

OpenStudy (anonymous):

I guess they wanted simple interest 15,500 x 0.025 x 8 = 3100.